A solution is made by mixing 6.5 g glucose with 26.7 g water, what is the percent by mass of glucose?

1
Expert's answer
2021-05-13T02:09:17-0400

"Mass\%(glucose)=\frac{m(glucose)}{m(glucose)+m(water)}\times100\%=\frac{6.5g}{6.5g+26.7g}\times100\%=19.6\%"


Answer: 19.6 %